i don't get how you do this work X^2+22x+21=0
Answer by checkley79(3341)

X^2+22x+21=0
(X+21)(X+1)=0
X+21=0
X=-21 ANS.
X+1=0
X=-1 ANS.
PROOF:
-21^2+22(-21)+21=0
441-462-21=0
441-441=0
0=0